### v1

# Layer by Layer: Uncovering Hidden Representations in Language Models

Abstract
From extracting features to generating text, the outputs of large language models (LLMs) typically rely on their final layers, following the conventional wisdom that earlier layers capture only low-level cues. However, our analysis shows that intermediate layers can encode even richer representations, often improving performance on a wide range of downstream tasks. To explain and quantify these hidden-layer properties, we propose a unified framework of representation quality metrics based on information theory, geometry, and invariance to input perturbations. Our framework highlights how each model layer balances information compression and signal preservation, revealing why mid-depth embeddings can exceed the last layer’s performance. Through extensive experiments on 32 text-embedding tasks and comparisons across model architectures (transformers, state-space models) and domains (language, vision), we demonstrate that intermediate layers consistently provide stronger features. These findings challenge the standard focus on final-layer embeddings and open new directions for model analysis and optimization, including strategic use of mid-layer representations for more robust and accurate AI systems.
Machine Learning, ICML
1 Introduction
Figure 1: Intermediate layers consistently outperform final layers on downstream tasks. The average score of 32 MTEB tasks using the outputs of every model layer as embeddings for three different model architectures. The x-axis is the depth percentage of the layer, rather than the layer number which varies across models.
Large Language Models (LLMs) have driven remarkable progress in natural language processing (NLP), achieving state-of-the-art results on many tasks (Brown et al., 2020; Devlin, 2018; Li et al., 2022). At the heart of most applications lies a common assumption: final-layer representations are the most useful for downstream tasks. Yet a fundamental question remains: does the final layer always yield the best representation?
In this paper, we conduct a layer-wise analysis of LLMs across diverse architectures—including Transformers (Vaswani et al., 2017), state-space models (SSMs) (Gu & Dao, 2024), and encoder-based models like BERT (Devlin, 2018)—spanning parameter scales from tens of millions to billions. Through systematic evaluation on 32 embedding tasks from the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2022), we find that intermediate layers often surpass the final layer by up to 16% in downstream accuracy. Figure 1 illustrates this phenomenon, where mid-depth layers provide particularly strong representations while the very last layer can become overly specialized to the pretraining objective.
A unified framework. To understand intermediate layers’ effectiveness, we integrate three complementary perspectives (Section 3):
  • Information-theoretic: How much do layers compress or preserve semantic information (Shwartz-Ziv & Tishby, 2019; Shwartz-Ziv, 2022)?
  • Geometric: How do token embeddings unfold in high-dimensional space (Hosseini & Fedorenko, 2023))?
  • Invariance: Are embeddings robust to input perturbations (e.g., InfoNCE (Oord et al., 2018), LiDAR (Thilak et al., 2024) and DiME (Skean et al., 2023))?
We show that these perspectives can be viewed under a single lens, which clarifies how intermediate layers strike a balance between retaining features and discarding noise.
Key findings and contributions. Our investigation leads to several important insights:
  • Intermediate layers consistently outperform final layers. This pattern is evident not only in Transformers but also in SSMs, suggesting a broad, architecture-agnostic effect.
  • Autoregressive vs. masked-language training. Autoregressive models exhibit a pronounced mid-layer “compression valley,” whereas masked or bidirectional models show milder intermediate changes.
  • Domain-general effect. We extend these results to vision models and find that autoregressive image transformers display the same mid-depth bottleneck, indicating that the training objective, rather than the data modality, is the key driver.
  • CoT finetuning. Analyzing chain-of-thought (CoT) reveals that finetuning can reshape mid-layer entropy, preserving latent context for multi-step reasoning.
Overall, our results challenge the default reliance on final-layer embeddings and highlight intermediate layers as potentially underutilized sources of meaningful features. In the rest of this paper, we detail our unified framework (Section 3), present an extensive set of experiments in both language and vision (Section 4, 6), and conclude with a discussion of implications for model design, training practices, and future directions.
2 Related Work
Understanding Neural Representations.
A long line of research has aimed to understand how deep neural networks encode and organize information. Early studies employed linear probes to interpret intermediate layers (Alain & Bengio, 2017), while subsequent efforts introduced more sophisticated techniques such as SVCCA (Raghu et al., 2017) to compare learned features across architectures and training regimes. Although these approaches shed light on representation dynamics, most focus on vision backbones or relatively shallow models. In contrast, our work extends layer-wise analysis to large-scale language models, highlighting specific behaviors of intermediate layers in autoregressive Transformers, state-space models (SSMs), and beyond.
Layer-wise Analysis in Language Models.
Transformer-based LLMs have sparked significant interest in which layers capture linguistic properties such as syntax and semantics (Liu et al., 2019; Tenney et al., 2019; Voita et al., 2019). More recent work (Jin et al., 2024; Gurnee & Tegmark, 2023; Fan et al., 2024) has shown that mid-depth layers sometimes hold surprisingly robust features, challenging the typical focus on final layers. Our contribution unifies and expands these observations through a large-scale, theoretical–empirical framework that quantifies the quality of every layer’s representation via information theory, geometry, and invariance metrics.
Architectural Comparisons.
Transformers remain the dominant architecture for NLP (Vaswani et al., 2017), but they come in multiple variants. Encoder-only models (e.g., BERT (Devlin, 2018)) typically use bidirectional attention and masked-language objectives, while decoder-only architectures (e.g., GPT (Brown et al., 2020)) follow an autoregressive paradigm. Meanwhile, newer state-space models (SSMs) such as Mamba (Gu & Dao, 2024) use recurrent-style dynamics for efficient long-sequence processing. Although these designs differ significantly in attention mechanisms and sequence modeling strategies, there has been little direct comparison of hidden-layer representations across them. In our work, we analyze Transformers (both encoder- and decoder-only) and SSMs under a common set of metrics, highlighting contrasts in how intermediate layers compress or preserve information and showing that intermediate-layer representations can excel across multiple architectures.
Representation Quality Metrics.
A variety of metrics have been proposed to quantify the “quality” of learned representations. We group them into three main categories:
  • Information-theoretic measures capture how much a model’s internal representations compress or preserve relevant information. For example, the Information Bottleneck (Shwartz-Ziv & Tishby, 2019; Shwartz-Ziv, 2022) analyzes whether intermediate layers discard noise while retaining essential features.
  • Geometric measures focus on the structure of embeddings in high-dimensional space. Classical approaches include analyzing singular values or effective rank of the representation matrix (Garrido et al., 2023), while more recent work explores curvature (Hosseini & Fedorenko, 2023) to quantify how smoothly tokens are mapped across consecutive positions or time steps.
  • Task-based or invariance metrics evaluate how well representations support downstream goals. For instance, augmentations-based approaches such as InfoNCE (Oord et al., 2018) and LiDAR (Thilak et al., 2024) estimate invariance to perturbations, while methods like NESum or Self-Cluster (Agrawal et al., 2022) link closely to entropy. In computer vision, these scores often correlate strongly with downstream accuracy, highlighting how robust the embeddings are.
Although these categories may appear distinct, we will show (Section 3) that many can be unified under a single lens. This unification illuminates why certain intermediate layers balance compression, geometry, and invariance so effectively, leading to better representations for downstream tasks.
Compression and Generalization.
Multiple lines of research connect compression and generalization performance (Deletang et al., 2024). For instance, Bordes et al. (2023) demonstrated that discarding certain layers in self-supervised encoders can even improve downstream accuracy, while Park et al. (2024a) found that LLM embeddings often lie in low-dimensional manifolds. Our empirical study reinforces these ideas by demonstrating that many networks—especially autoregressive Transformers—naturally develop a mid-layer bottleneck that appears crucial for balancing “signal” versus “noise.” We show how intermediate layers can achieve optimal trade-offs between preserving task-relevant information and discarding superfluous detail.
Overall, our work bridges these overlapping threads by evaluating a range of architectures and training paradigms via a unified set of metrics. Beyond merely confirming that intermediate layers can be effective, we elucidate why this happens, tying it to fundamental properties such as entropy, invariance, and geometry. This novel perspective provides an avenue for both finer-grained diagnostics of large language models and more deliberate design of mid-layer representations for downstream tasks.
3 A Unified Framework for Neural Representations
Key Takeaway: Matrix-based entropy unifies seemingly disparate metrics of representation quality, providing a single theoretical lens for analyzing compression, geometry, and invariance.
A central challenge in analyzing internal representations is determining how to assess their quality. Although existing work draws on numerous ideas—from mutual information to geometric manifold analysis to invariance under augmentations—these threads can seem disparate. In this section, we consolidate them into a unified theoretical framework that shows how these seemingly different metrics connect and why they collectively measure “representation quality.”
3.1 Notation and Motivation
Consider a neural network that maps inputs $\mathbf{x}$ (e.g., tokens in a sequence) to internal hidden states $\mathbf{Z}$. We denote $\mathbf{Z}\in\mathbb{R}^{N\times D}$ as a matrix of $N$ data samples (or tokens) in $D$ dimensions. Some key questions arise:
  1. 1.
    How compressed are these representations?
  2. 2.
    How robust are they to small perturbations or augmentations?
  3. 3.
    How do they geometrically organize different inputs?
The answers can illuminate which layers strike the right balance between preserving relevant features and discarding noise.
3.2 Matrix-Based Entropy: A Common Theoretical Thread
We focus on a key quantity known as matrix-based entropy (Giraldo et al., 2014; Skean et al., 2023), which applies directly to the Gram matrix $\mathbf{K}=\mathbf{Z}\mathbf{Z}^{\top}$. Let $\{\lambda_{i}(\mathbf{K})\}$ be the (nonnegative) eigenvalues of $\mathbf{K}$. For any order $\alpha>0$, define:
$S_{\alpha}(\mathbf{Z})\;=\;\frac{1}{1-\alpha}\,\log\!\biggl{(}\,\sum_{i=1}^{r}\!\Bigl{(}\tfrac{\lambda_{i}(\mathbf{K})}{\mathrm{tr}(\mathbf{K})}\Bigr{)}^{\alpha}\biggr{)},$ (1)
where $r=\mathrm{rank}(\mathbf{K})\leq\min(N,D)$. Intuitively, if only a few eigenvalues dominate, $S_{\alpha}(\mathbf{Z})$ is small—indicating a highly compressed representation. Conversely, if $\mathbf{Z}$ is spread out across many principal directions, $S_{\alpha}(\mathbf{Z})$ is large. By varying $\alpha$, one smoothly transitions between notions like collision entropy ($\alpha=2$) and von Neumann entropy ($\alpha\to 1$). We will typically use $\alpha=1$ for simplicity.
Implication: bridging geometry, invariance, and local vs. global features.
A key benefit of matrix-based entropy is that it unifies multiple representational perspectives:
  • Compression or information content: A handful of large eigenvalues in $\mathbf{K}=\mathbf{Z}\mathbf{Z}^{\top}$ indicates that $\mathbf{Z}$ is low-rank, i.e. the model has collapsed much of the input variation into fewer dimensions. In contrast, a more uniform eigenvalue spectrum implies higher-entropy, more diverse features.
  • Geometric smoothness: If tokens within a prompt follow a trajectory in embedding space with sharp turns, that curvature can manifest as skewed eigenvalue spectra (Hosseini & Fedorenko, 2023). Curvature also differentiates local transitions (token-to-token) from global structural patterns across longer segments or entire prompts.
  • Invariance under augmentations: Metrics like InfoNCE (Oord et al., 2018) and LiDAR (Thilak et al., 2024) effectively measure whether augmentations of the same sample (e.g. character swaps) map to similar embeddings. Strong invariance corresponds to stable clustering in $\mathbf{Z}\mathbf{Z}^{\top}$, which again depends on the distribution of eigenvalues and how local vs. global features are retained or discarded.
Thus, evaluating $S_{\alpha}(\mathbf{Z})$ provides a single lens for assessing “representation quality” across compression, geometric structure, and invariance—and highlights how both local details and global patterns are organized.
3.3 Representation Evaluation Metrics
Key Takeaway: Information-theoretic, geometric, and invariance-based metrics offer complementary perspectives on representation quality that can all be understood through matrix-based entropy.
We now introduce the seven representation evaluation metrics used in our experiments, grouped into three broad categories: (1) information-theoretic, (2) geometric, and (3) augmentation-invariance. All relate back to the Gram matrix $\mathbf{K}$ and hence to Eq. (1).
3.3.1 Information-Theoretic Metrics
Prompt Entropy.
Following Wei et al. (2024), we apply matrix-based entropy (Eq. 1) to the token embeddings within a single prompt. This prompt entropy quantifies how widely tokens are spread in the embedding space. Higher entropy indicates more diverse, less redundant token-level features; lower entropy implies stronger compression.
Dataset Entropy.
We can also aggregate embeddings across N prompts by taking the mean token embedding of each prompt to form $\overline{\mathbf{Z}}\in\mathbb{R}^{N\times D}$. Applying entropy to $\overline{\mathbf{Z}}$ yields a dataset-level measure of global diversity—revealing how distinctly the model separates different inputs.
Effective Rank  (Roy & Vetterli, 2007a)
can be shown to be a lower bound to $\exp(S_{1}(\mathbf{Z}))$, highlighting how dimensionality effectively shrinks if the representation is strongly compressed. We prove this connection later in Theorem 1. This has implications for popular representation evaluation metrics such as RankMe (Garrido et al., 2023) and LiDAR (Thilak et al., 2024), which are both inspired by Effective Rank.
3.3.2 Geometric Metrics
Curvature.
Proposed by Hosseini & Fedorenko (2023), curvature captures how sharply the token embeddings turn when viewed as a sequence in $\mathbb{R}^{D}$. For a prompt of length $L$, let $\mathbf{v}_{k}=\mathbf{z}_{k+1}-\mathbf{z}_{k}$ be the difference between consecutive tokens. The average curvature is:
$\bar{C}=\frac{1}{L-2}\sum_{k=1}^{L-2}\arccos\!\Bigl{(}\tfrac{\mathbf{v}_{k+1}^{\top}\mathbf{v}_{k}}{\|\mathbf{v}_{k+1}\|\|\mathbf{v}_{k}\|}\Bigr{)}.$
Higher curvature means consecutive tokens shift direction abruptly and more local level features; lower curvature suggests a smoother trajectory and global level features.
3.3.3 Augmentation Invariance Metrics
Lastly, we assess how stable the model’s representations are to small perturbations of the same input (e.g., random character swaps, keyboard-level changes; see Appendix). Suppose $p_{i}$ is augmented into $p_{i}^{(a)}$ and $p_{i}^{(b)}$. After embedding these, we compare the row vectors in $\mathbf{Z}_{1},\mathbf{Z}_{2}\in\mathbb{R}^{N\times D}$ under different scoring criteria:
InfoNCE.
This self-supervised objective (Oord et al., 2018) encourages matched samples to lie close in embedding space while pushing unmatched samples away. A lower InfoNCE loss indicates stronger invariance to augmentation.
LiDAR.
LiDAR (Thilak et al., 2024) uses a linear discriminant approach that measures within-class versus between-class scatter. Treating each prompt as its own class, LiDAR checks how well augmentations form tight clusters.
DiME.
Similarly, DiME (Skean et al., 2023) is grounded in matrix-based entropy. It compares real paired samples against random pairings to estimate how uniquely aligned correct augmentations are.
3.4 Core Theoretical Results
Key Takeaway: Our theoretical framework establishes concrete connections between representation entropy and downstream performance through properties like effective rank and invariance.
Here, we summarize key statements that justify why these metrics meaningfully measure representation quality. We refer to the appendix F for details and proofs. Beyond serving as a unifying view, matrix-based entropy also connects to foundational concepts like majorization, Schur concavity, and mutual information. Furthermore, we can directly relate the eigenvalue entropy to the matrix entropy, most naturally via the Effective Rank (Roy & Vetterli, 2007b). The following theorem makes this connection explicit.
Theorem 1 (Lower Bound via Effective Rank).
For Shannon-based entropy ($\alpha\to 1$),
$\mathrm{EffRank}(\mathbf{Z})\;\leq\;\exp\bigl{(}S_{1}(\mathbf{Z})\bigr{)},$
meaning a large effective rank implies a high entropy.
Under appropriate conditions on the data distribution and model, we can show connections between prompt entropy and dataset entropy via the following scaling behaviors:
Theorem 2 (Informal).
  1. 1.
    If prompt entropy remains near its maximum for all prompts, then the dataset entropy $S_{2}\!\bigl{(}\overline{\mathbf{Z}}\,\overline{\mathbf{Z}}^{\top}\bigr{)}$ grows on the order of $\log\!\bigl{(}\tfrac{D^{2}}{N}\bigr{)}.$
  2. 2.
    If prompt entropy instead stays near its minimum for all prompts, then dataset entropy grows more slowly, on the order of $\log\!\bigl{(}\tfrac{D^{2}}{N^{3}}\bigr{)}.$
In short, high token-level (prompt) diversity encourages broader global diversity in the dataset-level embeddings, whereas over-compressing token representations can limit how effectively different prompts separate. Our subsequent analysis connects these ideas to self-supervised objectives like InfoNCE, which also tie higher entropy to stronger robustness and discriminability in the learned representations.
Theorem 3 (Dataset Entropy Bounds InfoNCE).
For data $X$ and representation $Z(X)$, the InfoNCE loss on $N$ samples satisfies:
$\log(N)-\mathrm{InfoNCE}\;\;\leq\;\;I(X;Z)\;\;\leq\;\;H(Z),$
where $H(Z)$ is interpretable as matrix-based entropy at the dataset level. Hence, reducing InfoNCE implies learning a higher-entropy (and thus often more robust) representation.
Practical outlook.
Overall, our theoretical analysis shows that compression (entropy), geometry (curvature, rank), and invariance (e.g. InfoNCE) are all facets of how the Gram matrix $\mathbf{Z}\mathbf{Z}^{\top}$ distributes variance. Examining these metrics across different layers reveals exactly where a network “prunes” redundancy (low entropy) versus preserving essential distinctions (high entropy). This unified perspective also facilitates cross-architecture comparisons (e.g. Transformers vs. SSMs) by highlighting how each architecture organizes information internally. Beyond offering a theoretical foundation, it provides a practical blueprint for diagnosing, tuning, and improving hidden-layer representations.
(a) Prompt Entropy
(b) Curvature
(c) LiDAR
Figure 2: Pythia and Mamba’s intermediate layers show pronounced changes in representation quality metrics, while BERT’s remain more stable. Three representation evaluation metrics calculated on the wikitext dataset for every layer in Pythia-410M, Mamba 370M, and BERT-base architectures. The x-axis denotes layer depth as a percentage, allowing fair comparison between models with different layer counts.
4 Empirical Results
In this section, we empirically validate our theoretical framework through extensive experiments across architectures, scales, and training regimes. Our investigation centers on three key questions:
  • Do intermediate layers consistently outperform final layers across diverse downstream tasks?
  • How do these intermediate representations differ across architectures, training stages, and scales?
  • How does post-training methods (e.g., fine-tuning and chain-of-thought) reshape representations?
4.1 Downstream Task Performance
Key Takeaway: Intermediate layers of language models consistently outperform final layers across all architectures and tasks, challenging the conventional wisdom of using final-layer representations.
In this section, we use intermediate layers for downstream embedding tasks and employ our unified framework from Section 3, measuring all the embeddings across all layers.
4.1.1 Experimental Setup
Models
We evaluate three distinct architectural families: Pythia and Llama3 (decoder-only transformer) (Biderman et al., 2023; Dubey et al., 2024), Mamba (state space model) (Gu & Dao, 2024), BERT (encoder-only transformer) (Devlin, 2018) and LLM2Vec models (bidirectional attention) (BehnamGhader et al., 2024).
Tasks
We test each layer’s embeddings on 32 tasks from the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2022), spanning classification, clustering, and reranking. This comprehensive evaluation provides insight into how different layers capture task-relevant features. For a full list of the tasks, refer to the Appendix.
4.1.2 Intermediate Layers Often Outperform Final Layers
A key question is whether final-layer embeddings are indeed optimal for downstream tasks. In Figure 1, we compare average performance on MTEB tasks across all layers of the three models.
Key observation.
In nearly every task, some intermediate layer outperforms the final layer. The absolute improvement ranges from 2% to as high as 16% on average, and the best layer often resides around the mid-depth of the network. This phenomena is consistent across all the different architectures. This confirms emerging observations in recent work for generation tasks (Bordes et al., 2023; El-Nouby et al., 2024; Chen et al., 2020; Fan et al., 2024) and extends them to a wider range of benchmarks and tasks.
Why do these layers matter?
From our theoretical perspective, intermediate layers appear to strike a balance between retaining sufficient information (avoiding over-compression) and discarding low-level noise. Later in Section 4.2, we show that these sweet spots are not random but tied to how intermediate layers are processing information.
4.1.3 Layer-Wise Metrics Correlate with Downstream Performance
To validate our framework’s relevance, we analyze how each metric (entropy, InfoNCE, etc.) correlates with downstream performance. Figure 3 and Figure 8 show distance correlations between metrics and task scores for Pythia-410M. We make several key observations:
  • All metrics show strong relationships with performance
  • DiME, curvature, and InfoNCE exhibit particularly strong correlations
  • Associations remain robust across different correlation measures (Spearman, Kendall)
These relationships suggest that our metrics effectively capture what makes intermediate representations powerful for downstream tasks.
Figure 3: Relationship between representation metrics and task performance averaged across layers for Pythia 410M. Using distance correlation (dCor), we see strong associative relationships across the board with DiME exhibiting the strongest relationship with downstream performance. We use dCor due to its robustness and ability to measure both linear and non-linear relationships (dCor $\in[0,1]$ with 0 indicating statistical independence and 1 indicating strong dependency). We defer additional results to the Appendix.
4.2 Architectural and Scale Differences
Key Takeaway: Different architectures exhibit distinct patterns of information compression. Autoregressive models show mid-layer bottlenecks while bidirectional models maintain more uniform trends.
Aside from strong correlations with downstream performance, we can use our evaluation framework to assess the internal behaviors of LLMs. In both this section and Section 4.3, we use WikiText-103 (Merity et al., 2017) for analyzing our representation metrics on standard textual data. To investigate how architecture and model size influence representation quality, we compare three fundamentally different LLM variants—BERT (encoder-only), Pythia (decoder-only), and Mamba (state-space model)—and then scale up Pythia to observe emerging trends.
Encoder vs. Decoder vs. SSM.
Figure 2 shows how prompt entropy, curvature, and augmentation metrics evolve across each model’s layers. BERT, which encodes the entire input bidirectionally, generally maintains high entropy across layers, suggesting minimal compression: the model can see all tokens at once and need not discard as much information. By contrast, the decoder-only Pythia exhibits a strong mid-layer entropy dip, reflecting its autoregressive objective’s tendency to filter or prune non-local details in the middle of the network. As a result, Pythia’s “sweet spot” for downstream tasks often lies around mid-depth, where it balances essential context and compression. Mamba, meanwhile, processes sequences through a state-space approach that yields flatter, more uniform curves across depth: it neither retains as much information as BERT nor compresses as aggressively as Pythia’s mid-layers.
Scaling Size Effects.
In Figure 12, we analyze Pythia models ranging from 14M to 1B parameters. Larger Pythia models display more pronounced intermediate compression (entropy dips), indicating a heightened ability to distill relevant features. We also observe smoother token trajectories (lower curvature) and stronger invariance (e.g., higher LiDAR), consistent with findings that bigger models more effectively filter noise and capture long-range dependencies. Notably, these trends further reinforce why performance often peaks in the middle of the network: larger models gain more capacity to compress intermediate representations, yet still preserve crucial semantic details.
Finetuning Effects
In Figure 13, we study how finetuning affects the internal representations of Llama3 (Dubey et al., 2024). We compare the baseline Llama3-8B to two finetuned LLM2Vec models (BehnamGhader et al., 2024). The LLM2Vec-mntp-unsup-simcse model enables bidirectional attention in Llama3 and goes through two unsupervised training phases to improve Llama3’s performance on embedding tasks. The LLM2Vec-mntp-supervised adds an additional supervised finetuning phase. It is clear that both finetuned models have improved augmentation invariance. Furthermore, the unsupervised model has higher prompt entropy than Llama3 while the supervised model has less.
Layer-Level Analysis of Transformer Sub-Components.
While our experiments treat each Transformer layer as a single unit, Transformer blocks comprise multiple sub-layers (pre-attention normalization, self-attention, residuals, MLPs). By measuring entropy after each sub-layer, we find in Figure 15 that residual connections drive the mid-network compression observed in Section 4.2. Specifically:
  • Sub-layers before residuals (e.g. pre-attention, raw attention, or MLP pre-residual outputs) often show only mild compression; their representations still carry much of the original variability
  • Residual sub-layers exhibit a marked entropy drop, reflecting significant information filtering
The strong “valley” in entropy at intermediate layers is thus tied to how residual paths merge computed signals with the existing hidden state. This aligns with prior work indicating that residuals act as a regularizer or "noise filter" (Marion et al., 2024), smoothing out spurious components in hidden representations.
4.3 Impact of Training Progression
Takeaway: Significant changes during training occur in intermediate layers and early layers stabilize quickly, supporting the detokenization hypothesis.
We measure Pythia’s metrics at multiple checkpoints to understand how layer-wise representations evolve throughout training (Figure 4). Two main observations emerge:
(a) Prompt Entropy
(b) Curvature
(c) LiDAR
Figure 4: Strong trends in intermediate behavior emerge during training Representation evaluation metrics across layers at various Pythia-410M training checkpoints, ranging from step 1 to the final step at 143k. The x-axis is the model layer, showing how training affects different layers, while the colors are different checkpoints during training.
Intermediate Layers Undergo the Most Change.
The largest shifts in representation quality occur in mid-depth layers. Specifically, prompt entropy steadily decreases there as training progresses, implying that intermediate layers increasingly compress and abstract the input. Meanwhile, LiDAR scores are minimal in these same layers. Likewise, curvature becomes smoother in the middle of the network, suggesting the model refines its internal structure to capture longer-range or more nuanced patterns in language.
Early Layers Stabilize Quickly.
In contrast to the intermediate layers, the earliest layers change very little after the initial phase of training. This observation aligns with the “detokenization” hypothesis (Lad et al., 2024), which posits that early layers mainly convert raw tokens into a basic embedding space and then remain relatively fixed. As a result, the most substantial improvements in representation quality—such as enhanced compression—are driven primarily by the intermediate layers, reinforcing their importance for learning robust, high-level features.
4.4 Impact of Chain-of-Thought Finetuning
Key Takeaway: CoT finetuning enables models to maintain richer context throughout their layers.
Figure 5: Token-level prompt entropy across sequence length for Qwen 2.5 and Qwen 2.5-Math models. The base model (Qwen 2.5) exhibits greater prompt compression, while the finetuned model (Qwen 2.5-Math) maintains higher entropy, indicating more information retention.
Recent work has highlighted Chain-of-Thought (CoT) finetuning as a powerful strategy for improving reasoning capabilities (Arefin et al., 2024; DeepSeek-AI, 2025). To examine its effects, in Figure 5 we compare Qwen 2.5 and Qwen 2.5-Math (Yang et al., 2024a, b), where the latter underwent additional math pretraining and CoT finetuning. Measuring token-level prompt entropy across sequence length reveals that the finetuned model maintains higher entropy with lower variance across examples.
These findings suggest that CoT finetuning encourages models to preserve more context throughout their hidden layers, enabling better multi-step reasoning. Our framework provides a quantitative lens into how CoT fine-tuning pushes models to maintain richer internal representations across sequences, explaining its effectiveness in multi-step tasks. While CoT traces can be inspected directly in these models, our approach is particularly valuable for analyzing models that reason in continuous latent space (Hao et al., 2024).
5 Extreme Input Conditions
(a) Repetition
(b) Randomness
(c) Random Prompt Length
Figure 6: Prompt entropy across layers of Pythia 410M under various extreme input conditions. (a) Increasing token repetition leads to decreased entropy in intermediate layers. (b) Increasing token randomness results in higher entropy, especially in initial layers. (c) Unnormalized prompt entropy increases with prompt length due to the larger number of tokens. These results demonstrate how the model’s internal representations adapt to different types of input perturbations.
To gain a deeper understanding of the underline factors which effect the representation quality, we check how each layer responds to different types of inputs. We use Pythia 410M to three types of extreme prompts and measure prompt entropy across layers (Figure 6). We find that:
  1. 1.
    Token repetition compresses intermediate layers. As $p$ increases (i.e., more repeated tokens), prompt entropy decreases sharply in mid-depth layers. This indicates that the model effectively recognizes and encodes these repetitive patterns, discarding redundancy in its internal representation.
  2. 2.
    Random tokens inflate early-layer entropy. When we introduce token-level randomness, entropy increases significantly in the first few layers, revealing that these initial layers are especially sensitive to noise. By contrast, deeper layers appear more robust to such perturbations.
  3. 3.
    Prompt length raises raw entropy but grows sublinearly once normalized. Longer inputs naturally boost the unnormalized entropy because more tokens create more variation. However, normalized entropy expands at a slower rate, suggesting each additional token contributes less unique information.
Overall, these results confirm that intermediate layers play a major role in handling complex or unusual inputs, selectively compressing or filtering out repetitive patterns while retaining crucial distinctions. At the same time, early layers respond more sensitively to noise, and the incremental benefit of adding more tokens diminishes with prompt length. This behavior highlights the diverse ways in which different layers balance the trade-off between preserving and discarding information, further underscoring the unique strengths of intermediate representations.
6 Comparison to Vision Transformers
Although our focus has mainly been on language models, similar questions arise in computer vision. Vision architectures and training regimes differ widely, ranging from fully supervised methods to self-supervised approaches, and from bidirectional encoders to autoregressive transformers.
To investigate whether our findings generalize to vision models, we examine five representative vision approaches: ViT (Dosovitskiy et al., 2021), a supervised Transformer trained on labeled data; BEiT (Bao et al., 2022), a self-supervised encoder that reconstructs masked patches, analogous to masked token prediction in language; DINOv2 (Oquab et al., 2024), a self-supervised approach leveraging augmentations and exponential moving average teachers; MAE (He et al., 2022), a self-supervised framework that masks patches and reconstructs them, akin to masked autoencoders in language; and AIM (El-Nouby et al., 2024), an autoregressive Transformer that predicts the next patch in an image sequence (GPT-style next-token prediction). We evaluate each model on ImageNet-1k via layer-wise probing and our framework’s metrics.
A departure from language models for most vision Transformers.
Figure 14 shows that ViT, BEiT, DINOv2, and MAE exhibit strictly increasing downstream accuracy toward final layers, unlike language models. These models also show steadily increasing invariance metrics with depth, suggesting that without an autoregressive objective, vision Transformers have less need for drastic transformations at mid-depth.
AIM exhibits behavior similar to language models.
In contrast, AIM—which is explicitly autoregressive over image patches—shows an entropy “valley” and corresponding peak in downstream accuracy at its intermediate layers (El-Nouby et al., 2024). This mimics the patterns we observe in LLMs like Pythia, suggesting that autoregressive training induces an information bottleneck mid-depth. As in language modeling, forcing a strictly left-to-right (or patch-to-patch) prediction can drive the model to compress non-local details earlier, then re-expand relevant features.
Autoregression as the driving factor.
Taken together, these results indicate that the strong mid-layer compression observed in LLMs is not purely a property of “sequential token data” vs. “image patch data,” but rather a byproduct of autoregressive training. While various self-supervised (or fully supervised) objectives in vision often foster more uniform feature building across layers, autoregressive vision models develop the same mid-layer bottlenecks and sweet spots that we see in language. Thus, the architectural and objective design—especially whether or not a model is autoregressive—appears crucial in shaping layer-wise representation quality, regardless of domain.
7 Discussion and Conclusion
In this work, we investigate the representation quality of intermediate layers in LLMs, shedding light on their critical role in downstream task performance. We introduce a unified framework of evaluation metrics, establish theoretical connections among them, and apply these metrics to analyze Transformer-based architectures, SSMs, and vision models. One key phenomenon unveiled by prompt entropy was an information bottleneck in the middle layers of autoregressive transformers in both vision and language domains. Furthermore, our results reveal that intermediate layers often surpass final layers in representation quality, emphasizing their importance for feature extraction. DiME, curvature, and infoNCE correlate very well with downstream performance, suggesting a fundamental connection between representation and generalizability.
In conclusion, our study deepens the understanding of internal representation dynamics in LLMs. These insights not only enrich the theoretical foundations of model representations but also offer practical implications for optimizing model design, training strategies, and real-world applications. Future research could investigate the underlying causes of intermediate layer compression and develop specialized metrics tailored to LLMs, enabling more precise and effective representation evaluation.
Impact Statement
Our paper studies the inner workings of large language models with findings that may challenge typical assumptions about the importance of intermediate layers in large language models and the representations they learn. Our findings suggest that representations from these layers can yield better performance on a variety of downstream tasks, which can have implications for model interpretability, robustness, and efficiency.
From an ethical standpoint, the ability to leverage intermediate-layer representations could impact fairness and bias considerations in evaluating model performance or in model deployment. By helping better identify latent features and representations, our approach may amplify latent biases. We welcome and encourage future work to explore methods that can ensure that intermediate-layer representations do not disproportionately reinforce biases or lead to unintended disparities in real-world applications.
8 Acknowledgements
Oscar Skean is supported by the Office of the Under Secretary of Defense for Research and Engineering under award number FA9550-21-1-0227.
Appendix A Architectural Details
In this section, we elaborate on the specific architectures of Transformers and State Space Models (SSMs). We outline the mathematical foundations, including the weight matrices, attention mechanisms for Transformers, and the state transition matrices for SSMs. Detailed equations and parameter configurations are provided to facilitate replication and deeper understanding.
A.1 Transformer
The Transformer architecture (Vaswani et al., 2017) utilizes self-attention mechanisms. Given an input $\mathbf{x}$, the key ($\mathbf{K}$), query ($\mathbf{Q}$), and value ($\mathbf{V}$) matrices are computed as:
$\mathbf{Q}=\mathbf{x}\mathbf{W}_{Q},\quad\mathbf{K}=\mathbf{x}\mathbf{W}_{K},\quad\mathbf{V}=\mathbf{x}\mathbf{W}_{V},$ (2)
where $\mathbf{W}_{Q},\mathbf{W}_{K}\in\mathbb{R}^{d\times d_{k}}$ and $\mathbf{W}_{V}\in\mathbb{R}^{d\times d_{v}}$ are learned weights.
The attention weights are calculated using:
$\mathbf{A}=\operatorname{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}+\mathbf{M}\right),$ (3)
where $\mathbf{M}$ is a mask to enforce causality in autoregressive tasks.
The output is then:
$\mathbf{y}=\mathbf{A}\mathbf{V}.$ (4)
A.2 State Space Models
SSMs (Gu & Dao, 2024) model sequences using recurrent dynamics. The hidden state $\mathbf{h}_{t}$ and output $\mathbf{y}_{t}$ at time $t$ are updated as:
$\mathbf{h}_{t}$ $=\mathbf{A}\mathbf{h}_{t-1}+\mathbf{B}\mathbf{x}_{t},$ (5)
$\mathbf{y}_{t}$ $=\mathbf{C}\mathbf{h}_{t}+\mathbf{D}\mathbf{x}_{t},$ (6)
where $\mathbf{A}\in\mathbb{R}^{n\times n}$, $\mathbf{B}\in\mathbb{R}^{n\times d}$, $\mathbf{C}\in\mathbb{R}^{d\times n}$, and $\mathbf{D}\in\mathbb{R}^{d\times d}$ are learned parameters.
Appendix B Discussion on Prompt Entropy
Figure 7: The behavior of Eq. 1 for varying values of $\alpha$ on Gram matrices with eigenvalues distributed with a $\beta$-power law such that $\lambda_{i}=i^{-\beta}$.
The first measure of token embedding diversity we call prompt entropy. This entropy is measured on the intermediate tokens and captures how diverse the token representations are.
We follow the work of (Wei et al., 2024) and use $\alpha$-order matrix-based entropy (Giraldo et al., 2014; Skean et al., 2023, 2024), which serves as a tractable surrogate for traditional Rényi’s $\alpha$-order entropy (Rényi, 1961). The quantity is calculated using a similarity kernel $\kappa$ on a batch of samples drawn from a distribution, without making explicit assumptions on what the true distribution is. The choice of kernel $\kappa$ is flexible and can be any infinitely divisible kernel such as the Gaussian kernel, linear kernel, or Laplacian kernel, among others. For this work, we restrict ourselves to the linear kernel $\kappa(a,b)=ab^{T}$. This choice is motivated by the linear representation hypothesis (Park et al., 2024b) which finds that large language model representations encode high-level concepts such as truth (Burns et al., 2022), honesty (Mallen & Belrose, 2024), and part-of-speech (Mamou et al., 2020) in linearly separable manifolds.
The equation for matrix-based entropy was previously defined in Eq. 1. One way to interpret Eq. 1 is as the $\alpha$-order Rényi entropy of the Gram matrix eigenvalues [^1] The non-zero eigenvalues of the Gram matrix $ZZ^{T}$ are equivalent to those of the covariance matrix $Z^{T}Z$. Using the covariance matrix instead of the Gram matrix in Eq. 1 makes no difference and is more computationally efficient if $D<N$.. Notice how each eigenvalue is divided by $\textrm{tr}(\mathbf{K}_{\mathbf{Z}})$ before being raised to the $\alpha$ power. This is so that the eigenvalues of $\mathbf{K}_{\mathbf{Z}}$ sum to one (because $\textrm{tr}(\cdot)=\sum_{i=1}^{n}\lambda_{i}(\cdot)$), which is a necessary condition to treat the eigenvalues as a probability distribution. Futhermore, each eigenvalue of $\mathbf{K}_{\mathbf{Z}}$ signifies the variance of samples in a particular principal component direction (Scholkopf & Smola, 2018). If entropy is low, then the eigenvalues form a heavy-tail distribution which implies that a few components dominate the variance of samples in $Z$. On the other hand, at maximum entropy, the eigenvalues form a uniform distribution and samples are spread equally in all directions. Matrix-based entropy is reminiscent of the LogDet entropy which uses the determinant of $\mathbf{K}_{\mathbf{Z}}$ to capture how much "volume" a dataset occupies (Shwartz-Ziv et al., 2023; Zhouyin & Liu, 2021). The LogDet entropy is given by $S_{\textrm{LogDet}}(Z)=\log\det(\mathbf{K}_{\mathbf{Z}})-\log 2$. One can use Jensen’s inequality to show that the LogDet entropy is a lower bound of Eq 1 when $\lim_{\alpha\rightarrow 1}$ (Appendix J.4 of (Shwartz-Ziv et al., 2023)).
Depending on the choice of $\alpha$, several special cases of matrix-based entropy can be recovered. In particular, when $\lim_{\alpha\rightarrow 1}$ it equals Shannon entropy (also referred to as von Neumann entropy in quantum information theory (Bach, 2022; Boes et al., 2019)), and when $\alpha=2$ it equals collision entropy. Interestingly, the case of $\alpha=2$ can be calculated without explicit eigendecomposition (Skean et al., 2024). We show in the Appendix Figure 7 how varying values of $\alpha$ affect the matrix-based entropy of Gram matrices with eigenvalues distributed with a $\beta$-power law such that $\lambda_{i}=i^{-\beta}$. It is shown that for larger values of $\alpha$, smaller eigenvalues contribute more to the entropy.
Appendix C Dataset Details
C.1 Wikitext Dataset
We used the wikitext dataset (Merity et al., 2017) for the majority of our experiments in Sections 4.2 and 5. This was downloaded from Salesforce/wikitext on huggingface. The dataset consists of 100 million tokens scraped from the Featured articles on wikipedia. We filtered out prompts which were less than 30 tokens or were wikipedia section headings.
C.2 MTEB
The 32 tasks we used from the Massive Text Embedding Benchmark (MTEB) are detailed in Table 1. They are English language tasks covering clustering, classification, reranking, and sentence-to-sentence.
Task Domain Tasks # Tasks (32 Total)
Pair Classification SprintDuplicateQuestions, TwitterSemEval2015, TwitterURLCorpus 3
Classification AmazonCounterfactualClassification, AmazonReviewsClassification, Banking77Classification, EmotionClassification, MTOPDomainClassification, MTOPIntentClassification, MassiveIntentClassification, MassiveScenarioClassification, ToxicConversationsClassification, TweetSentimentExtractionClassification 10
Clustering ArxivClusteringS2S, BiorxivClusteringS2S, MedrxivClusteringS2S, RedditClustering, StackExchangeClustering, TwentyNewsgroupsClustering 6
Reranking AskUbuntuDupQuestions, MindSmallReranking, SciDocsRR, StackOverflowDupQuestions 4
Sentence to Sentence BIOSSES, SICK-R, STS12, STS13, STS14, STS15, STS16, STS17, STSBenchmark 9
Table 1: MTEB Tasks used in experiments covering a wide range of different use-cases and domains.
Appendix D Prompt Augmentations
For the augmentation-invariance metrics such as infoNCE, LiDAR, and DiME, we use the NLPAug library (Ma, 2019) to augment our prompts. We use three types of augmentations.
  • The SplitAug augmentation randomly splits words into two parts by adding a space.
  • The RandomCharAug augmentation randomly inserts, substitutes, swaps, or deletes characters.
  • The Keyboard augmentation randomly substitutes characters with other characters that are at a distance of one as measured on a QWERTY keyboard. For instance, the character "k" may be replaced with "i", "l", "m", or "j".
We use the pseudocode below to do our augmentations using three types of augmentations, using the default library settings for each type. When computing augmentation-invariance metrics like infoNCE or DiME, we use the two augmented prompts rather than using one augmented prompt alongside the original prompt. Note that these augmentations may change the token length $T$ of a prompt.
aug = naf.Sequential([
naw.SplitAug(p=0.3),
nac.RandomCharAug(p=0.3),
nac.KeyboardAug(p=0.3),
(aug_A, aug_B) = aug.augment(prompt, num_augmentations=2)
prompt -> "The quick brown fox jumps over the lazy dog."
aug_A -> "The quDUk b rown fox wEmps o ver the l azy dog."
aug_B -> "The qTuXi bro wn fox uVm)s ob3r the la_k dog."
Appendix E Extreme Prompts
E.1 Increasing Repetition
We take regular prompts from the wikitext dataset, tokenize them, and then for each token we randomly replace it with probability $p$. We draw replacements tokens by sampling a random token from within the prompt. We show examples below for varying levels of $p$.
  • ($p=0$)   Mint records indicate the first gold dollars were produced on May 7…
  • ($p=0.1$) Mint records indicate the first gold dollars were Mint Mint May 7…
  • ($p=0.5$) Mint records Mint Mint Mint gold dollars were Mint Mint Mint 7…
  • ($p=1.0$) Mint Mint Mint Mint Mint Mint Mint Mint Mint Mint Mint Mint Mint…
E.2 Increasing Randomness
We take regular prompts from the wikitext dataset, tokenize them, and then for each token we randomly replace it with probability $p$. We draw replacements uniformly from the tokenizer distribution. We show examples below for varying levels of $p$. Unlike the character-level random noise added to prompts in Section with random noise discussed in Appendix D which might change the number of tokens $T$ of the prompt, the token-level random noise used here does not do so.
  • ($p=0$)   Mint records indicate the first gold dollars were produced on May 7…
  • ($p=0.1$) Mint records indicate salivary first gold dollars were produced on May NaCl…
  • ($p=0.5$) Mint records Dallas actively first dollars persufors on Mayder129 18…
  • ($p=1.0$) arf emulsion minorensteinorianmega_TOStack potsRecip Installifykeeping…
Appendix F Theorems
Definition 1.
(Majorization) Let $p,q\in\mathbb{}{R}^{n}$ be nonnegative vectors such that $\sum_{i=1}^{N}p_{i}=\sum_{i=1}^{N}q_{i}$. We say that q majorizes p, denoted by $p\preccurlyeq q$, if their ordered sequences $p_{[1]}\geq\cdots\geq p_{[n]}$ and $q_{[1]}\geq\cdots\geq q_{[n]}$ satisfy:
$\sum_{i=1}^{k}p_{[i]}\leq\sum_{i=1}^{k}q_{[i]}\textrm{\quad for \quad}k=1,\cdots,n$ (7)
Definition 2.
(Schur-Convexity) A real-valued function $f$ on $\mathbb{R}^{n}$ is called Schur-convex if $p\preccurlyeq q\implies f(p)\leq f(q)$, and Schur-concave if $p\preccurlyeq q\implies f(q)\leq f(p)$.
Lemma 1.
The matrix-based entropy, as given in Equation 1, is a Schur-concave function for $\alpha>0$. This result is well-known and, for instance, was recently given by Lemma 4.1 in (Giraldo et al., 2014).
Theorem 4.
Suppose we have a matrix of embeddings $Z\in\mathbb{R}^{N\times D}$ and its covariance $Z^{T}Z$. Then the effective rank of $Z$ is an lower bound of $\exp(S_{1}(Z))$, where $S_{1}$ denotes the matrix-based entropy of $\alpha=1$.
Proof.
Denote the ordered singular values of $Z$ as $\sigma_{1}\geq\cdots\geq\sigma_{\min{(N,D)}}\geq 0$ and the ordered eigenvalues of $Z^{T}Z$ as $\lambda_{1}\geq\cdots\geq\lambda_{\min{(N,D)}}\geq 0$. Without loss of generality, assume that $\sum_{i=1}^{N}\sigma_{i}=\sum_{i=1}^{N}\lambda_{i}=1$. If this is not the case, then set $\sigma_{i}\coloneq\frac{\sigma_{i}}{\sum_{i=1}^{N}\sigma_{i}}$ and $\lambda_{i}\coloneq\frac{\lambda_{i}}{\sum_{i=1}^{N}\lambda_{i}}$.
It is straightforward to show that $\sigma_{i}^{2}=\lambda_{i}$. Because $\forall i\quad\sigma_{i}\leq 1$, we have that $\sigma_{i}\geq\lambda_{i}$. This implies that $\lambda\preccurlyeq\sigma$. Therefore, $S_{1}(\sigma)\leq S_{1}{(\lambda)}\implies\textrm{effective rank}(Z)\leq\exp{S_{1}{(Z)}}$.
Proposition 1.
(Random Unit Vectors are Nearly Orthogonal) Suppose we have $m$ unit vectors in $\mathbb{R}^{d}$, that are distributed according to the uniform distribution on the hyper-sphere. Then with probability at least $1-m^{2}\sqrt{2\pi}e^{\frac{-n\epsilon^{2}}{2}}$, we have that for any pair $i,j$, $i\not=j$,
$\langle\mathbf{v_{i}},\mathbf{v_{j}}\rangle\leq\epsilon.$
Proof.
Notice that the probability of not landing in the $\epsilon$ band $T_{\epsilon}\subset\mathbb{S}_{n-1}$ around the equator of the hypersphere of dimension $n$ can be bounded as,
$\mathbb{P}(T_{\epsilon}^{c})\leq\sqrt{2\pi}e^{\frac{-n\epsilon^{2}}{2}}.$
Which is a result from (Wainwright, 2019). Notice that for sufficiently small epsilon, this means that the dot product of any two, randomly chosen vectors can be bounded with high probability. Indeed notice that,
$\arccos(\langle\mathbf{v_{i}},\mathbf{v_{j}}\rangle)=\theta,$
and that if $\mathbf{v_{j}}\in T_{\epsilon}$, then, treating $\mathbf{v_{i}}$ as $e_{1}$, the basis vector, without loss of generality, we have that,
$\theta\leq\arccos(\frac{\epsilon}{2}).$
So then,
$\langle\mathbf{v_{i}},\mathbf{v_{j}}\rangle\leq\cos(\arccos(\frac{\epsilon}{2}))=\frac{\epsilon}{2}\leq\epsilon.$
Now, by the union bound on each $i\not=j$, we get that,
$\mathbb{P}(\exists i,j:\langle\mathbf{v_{i}},\mathbf{v_{j}}\rangle>\epsilon)$ $\leq\sum_{i\not=j}\mathbb{P}(\langle\mathbf{v_{i}},\mathbf{v_{j}}\rangle>\epsilon)$
$\leq m^{2}\sqrt{2\pi}e^{\frac{-n\epsilon^{2}}{2}}.$
So then with probability at least $1-m^{2}\sqrt{2\pi}e^{\frac{-n\epsilon^{2}}{2}}$, we have that, for any pair $i,j$,
$\langle\mathbf{v_{i}},\mathbf{v_{j}}\rangle\leq\epsilon.$
Theorem 5.
(Maximum Prompt Entropy implies Large Dataset Entropy.) Suppose we have a orthogonally equivarient representation model $Z$ such that for all sequences $Z_{i}=Z(X_{i})$ the prompt entropy is maximal and the rows are unit. Suppose also that the data distribution $\mathbf{Data}$ is a isotropic unit Gaussian. Suppose we draw sequences of length $L=D$ from the data distribution. Then with probability $1-N^{2}\sqrt{2\pi}e^{\frac{-n\epsilon^{2}}{2N^{2}}}$ over draw of $\{\mathbf{x_{i}}\}_{i=1}^{N}\sim\mathbf{Data}$, we have that,
$|e^{-S_{2}(QQ^{\top})}-\frac{N}{D^{2}}|\leq\epsilon$
Proof.
First note that, since the prompt entropy is maximal for each sample $Z(X_{i})$, then the matrix $K_{Z}=ZZ^{\top}$ is full rank. Since by assumption each row of $Z$ has unit rows, then we know that $\|Z\|_{F}^{2}=L=\sum_{k=1}^{L}\sigma_{k}^{2}$. In particular we also know that $\sigma_{i}=\sigma_{j}$ for all pairs $i,j$ by the assumption that the prompt entropy is maximized. In particular we then know that $ZZ^{\top}$ is a orthogonal matrix, and the rows of $Z$ form an orthonormal set. We can then write, for some $O_{i}$ a rotation matrix, that,
$\mathbf{q_{i}}=\frac{1}{D}\sum_{i=1}^{D}\mathbf{z_{i}}=\frac{1}{D}O_{i}\mathbf{1}.$
Since by assumption our model $Z(\cdot)$ is orthogonally equivarient, and the Data distribution is radially symmetric, it follows that these $\{\mathbf{q_{i}}\}_{i=1}^{N}$ are random points on the hypersphere of radius $\frac{1}{\sqrt{D}}$. This means that the matrix $\sqrt{D}Q$ consists of rows that are uniform points on hypersphere of radius $1$. Now notice that,
$\|QQ^{\top}\|_{F}^{2}$ $=\frac{1}{D^{2}}\|DQQ^{\top}\|_{F}^{2}$
=$\frac{1}{D^{2}}(\sum_{i=1}^{N}\|\sqrt{D}q_{i}\|^{2}+\sum_{i\not=j}\langle\sqrt{D}q_{i},\sqrt{D}q_{j}\rangle).$
Since $\sqrt{D}q_{i}$ is a unit vector this will simplify to,
$\|QQ^{\top}\|_{F}^{2}=\frac{1}{D^{2}}(N+\sum_{i\not=j}\langle\sqrt{D}q_{i},\sqrt{D}q_{j}\rangle).$
Now notice that by proposition, we have that with probability at least $1-N^{2}\sqrt{2\pi}e^{\frac{-D\epsilon^{2}}{2N^{2}}}$,
$\forall i\not=j:\langle\mathbf{v_{i}},\mathbf{v_{j}}\rangle\leq\frac{\epsilon}{N}.$
The union bound then tells us that,
$\mathbb{P}(\forall i\not=j:|\langle\sqrt{D}q_{i},\sqrt{D}q_{j}\rangle|\leq\frac{\epsilon}{N^{2}})\geq 1-N^{2}\sqrt{2\pi}e^{\frac{-D\epsilon^{2}}{2N^{2}}}.$
So then with probability at least $1-N^{2}\sqrt{2\pi}e^{\frac{-D\epsilon^{2}}{2N^{2}}}$ over the draw of the data points, we have that,
$\ |\|QQ^{\top}\|_{F}^{2}\ -\ \frac{N}{D^{2}}|\leq\epsilon.$
So then since,
$S_{2}(QQ^{\top})=\log\left(\frac{1}{\|QQ^{\top}\|_{F}^{2}}\right),$
we have that, $e^{-S_{2}(QQ^{\top})}=\|QQ^{\top}\|_{F}^{2}$. In particular,
$|e^{-S_{2}(QQ^{\top})}-\frac{N}{D^{2}}|\leq\epsilon.$
Which completes the proof. ∎
Theorem 6.
(Minimal Prompt Entropy Implies Small Dataset Entropy 2) Suppose we have a orthogonally equivarient representation model $Z$ such that for all sequences $Z_{i}=Z(X_{i})$ the prompt entropy is minimal and the rows are unit. Suppose also that the data distribution $\mathbf{Data}$ is a isotropic unit Gaussian. Suppose we draw sequences from the data distribution. Then with probability $1-N^{2}\sqrt{2\pi}e^{\frac{-D^{3}\epsilon^{2}}{2N^{6}}}$ over the draw of $\{\mathbf{x_{i}}\}_{i=1}^{N}\sim\mathbf{Data}$, we have that,
$|e^{-S_{2}(QQ^{\top})}-\frac{m^{3}}{N^{2}}|\leq\epsilon.$
Proof.
Since the prompt entropy is minimal for each sample, we know that each $Z(X_{i})$ will be a rank one matrix, so we can write it as the outer product. In particular, we can write $Z(X_{i})=\mathbf{v_{i}}{\mathbf{u_{i}}}^{\top}$. However, since the rows of $Z(X_{i})$ are of unit length, we know that all the rows are identical, so we may write $Z(X_{i})=\mathbf{v_{i}}\mathbf{1}^{\top}$. Then, it follows that,
$\mathbf{q_{i}}=\frac{1}{D}\sum_{i=j}^{D}\mathbf{z_{j}^{i}}=\frac{N}{D}\mathbf{v}_{i}.$
In particular the matrix $\frac{D}{N}Q$ has rows that are all unit vectors, and these are randomly distributed uniformly on the hypersphere. Now notice that,
$\|QQ^{\top}\|_{F}^{2}$ =$\sum_{i=1}^{N}\|q_{i}\|^{2}+\sum_{i\not=j}\langle q_{i},q_{j}\rangle$
=$\sum_{i=1}^{N}\frac{N^{2}}{D^{2}}\|v_{i}\|^{2}+\sum_{i\not=j}\frac{N^{2}}{D^{2}}\langle v_{i},v_{j}\rangle$
=$\frac{N^{3}}{D^{2}}+\sum_{i\not=j}\frac{N^{2}}{D^{2}}\langle v_{i},v_{j}\rangle.$
Now by proposition, with probability at least $1-N^{2}\sqrt{2\pi}e^{\frac{-D^{3}\epsilon^{2}}{2N^{6}}}$, we know that, for all $i\not=j$,
$\langle\mathbf{v_{i}},\mathbf{v_{j}}\rangle\leq\frac{\epsilon D^{2}}{N^{3}}.$
So then,
$|\|QQ^{\top}\|_{F}^{2}-\frac{N^{3}}{D^{2}}|\leq\epsilon.$
In particular,
$|e^{-S_{2}(QQ^{\top})}-\frac{N^{3}}{D^{2}}|\leq\epsilon.$
Theorem 7.
(Dataset Entropy Bounds InfoNCE) Let $X\sim\textbf{Data}$ be a discrete random variable distributed according to the data distribution. Let $X\to Z$ be the Markovian relation between $X$ and the representation $Z$. Then, the InfoNCE loss on $N$ samples from Data satisfies,
$\log(N)-\text{InfoNCE}\leq I(X;Z)\leq H(Z).$
The entropy $H(Z)$ is analogous to the Dataset Entropy.
Proof.
The first inequality follows as a simple result from (Oord et al., 2018). Then, use that,
$I(X;Z)=H(Z)-H(Z|X)\leq H(Z).$
Appendix G Additional Plots & Visualizations
Figure 8: Relationships between representation metrics and task performance averaged across layers for Pythia 410M. Using a variety of linear and non-linear measures—Spearman’s $\rho$, Kendall’s $\tau$, and distance correlation (dCor)—we see strong inversely associative relationships with the exception of InfoNCE which shows a positive, but still strong associativity. Ranges of $\rho,\tau\in[-1,1]$ and dCor $\in[0,1]$ with 0 indicating independence and 1 indicating strong dependency.
Figure 9: Relationship between representation metrics and task performance averaged across layers for BERT. Using distance correlation (dCor), we see strong associative relationships across the board with LiDAR and dataset entropy exhibiting the strongest relationship with downstream performance. We use dcor due to its robustness and ability to measure both linear and non-linear relationships (dCor $\in[0,1]$ with 0 indicating statistical independence and 1 indicating strong dependency). Other correlative measures also indicate moderate to strong relationships.
(a) Prompt Entropy
(b) Dataset Entropy
(c) Curvature
(d) infoNCE
(e) LiDAR
(f) DiME
Figure 10: textbfPythia’s intermediate layers show pronounced changes in representation quality metrics, while Mamba’s remain more stable. Representation evaluation metrics across layers in Pythia 410M and Mamba 370M architectures. The x-axis denotes model depth as a percentage, allowing fair comparison between models with different layer counts.
Figure 11: Representation evaluation metrics across layers at various training checkpoints, ranging from step 1 to the final step at 143k. The x-axis represents the depth percentage of the model, showing how training affects different layers, particularly in the intermediate stages.
(a) Prompt Entropy
(b) Curvature
(c) LiDAR
Figure 12: Pythia and Mamba’s intermediate layers show pronounced changes in representation quality metrics, while BERT’s remain more stable. Three representation evaluation metrics calculated on the wikitext dataset for every layer in Pythia-410M, Mamba 370M, and BERT-base architectures. The x-axis denotes layer depth as a percentage, allowing fair comparison between models with different layer counts.
(a) Prompt Entropy
(b) Curvature
Figure 13: Finetuning affects the internal behavior of LLMs. Representation evaluation metrics across layers for Llama3 and two finetuned versions of Llama3.
Figure 14: Comparison of vision transformers trained with different pretext tasks.
Figure 15: Behavior of effeective rank at different stages within a transformer block.